OpenStudy (anonymous):

find the period,range and amplitude of the cosine function y=3cos θ.

OpenStudy (phi):

match with \[ A \text{ cos}(\frac{2 \pi}{T} \theta + \phi ) \] where A is the amplitude T is the period \( \phi \) is the phase

OpenStudy (phi):

the range is the y-values : the numbers from the min to the max.